OpenStudy (anonymous):

Solve using the multiplication principle - 1/6x= -9/10

jimthompson5910 (jim_thompson5910):

- 1/6x= -9/10 -6(- 1/6x)= (-9/10)*(-6) x = 54/10 x = 27/5

OpenStudy (saifoo.khan):

\[-\frac 16x = -\frac {9}{10}\]\[\frac 16x = \frac {9}{10}\]Cross multiply. \[x = \frac{9}{10} \times 6\]

OpenStudy (saifoo.khan):

\[x = \frac{54}{10}\]

OpenStudy (anonymous):

thanks guys

OpenStudy (saifoo.khan):

x = 27/5